OpenStudy (anonymous):

Mr. Davis has 15 students. He wants to send 3 of his students to pick up books. How many combinations of students can he choose.

OpenStudy (anonymous):

15! i think

OpenStudy (anonymous):

15*14*3*12*11*10*9*8*7*6*5*4*3*2*1

OpenStudy (cwrw238):

15C3 - that number of combinations of 3 out of 15 work it out as follows 15 * 14 * 13 ---------- 3 * 2 * 1

OpenStudy (cwrw238):

what you have there tyler is number of possible arrangements of 15

OpenStudy (anonymous):

oh ok im working on the same thing so i have to practice a little more

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

455?

OpenStudy (cwrw238):

5 * 7 * 13 = 13*35 = 455 - yup